What is the change in energy of a system that requires 80 J of energy to do 150 J of work?


1
Expert's answer
2022-02-16T11:25:01-0500

According to the 1st Law of Thermodynamics,ΔQ= ΔU + ΔW , i.e., change is heat exchange will be equal to change in internal energy and the work done.

Now coming to the question, considering the gas as system.

Given that heat is supplied = +80 cal = +(80 x 4.18) Joule = +334.4 Joule

and the work done by the gas = +150 joule.

Now applying the 1st law, we get,

ΔU = ΔQ - ΔW = (334.6 - 150) J= 184.4 Joule or (184.4/4.14) cal = 44.114 cal